# This version: August 29, 2025 )

### v3

@@ 5.2.2 Treatment Effects
Figure 4 presents the estimated bounds. The left panel shows treatment effects on the probability of graduating from Medicine at PUC Santiago; the right panel shows treatment effects on the probability of graduating from a top university. Starting with the left panel, we find, reassuringly, that assignment to Medicine at PUC increases the probability of graduating from that program. In all five cases, the bounds derived under WPO+UMAS+fields exclude zero, identifying a positive treatment effect. Importantly, the magnitude of the effect varies across next-best alternatives. For example, students whose second-best is Science at PUC appear to benefit less from admission to Medicine at PUC than those whose next-best is Medicine at UChile or at Universidad de Santiago de Chile. These differences are reflected in non-overlapping bounds across several cases.
The right panel reveals similar patterns. Admission to Medicine at PUC increases the probability of graduating from a top university in four of the five cases, and regardless of whether the next-best option is at a top university itself. For instance, the estimated bounds indicate a positive effect both for students whose next-best program is Science or Engineering at PUC and for those whose next-best is Medicine at Universidad de Concepción or Universidad de Santiago de Chile. There again, the bounds under the WPO+UMAS+fields assumption set highlight substantial heterogeneity across next-best alternatives.
As we relax assumptions, the bounds widen and become less informative—their ability to recover the sign of the treatment effect diminishes. This exercise thus helps clarify the identifying power of each assumption, providing researchers with a framework to evaluate the empirical content of different preference assumptions in similar settings.
@@
The figure shows estimated bounds for the average effect on the probability of graduation (y-axis) of assigning students to Medicine at PUC as opposed to assigning them to a second-best program. Bounds are shown for five different second-best programs (x-axis), two binary outcomes (left and right panels), and four behavioral assumptions used to construct $\bm{\mathrm{Q}}_{j}$ (four bars for each second-best program). The outcomes are graduation from Medicine at PUC on the left panel and graduation from a top university (PUC or UChile) on the right panel. The four behavioral assumptions are, from left to right, SPO, WPO in combination with UMAS and fields, WPO in combination with UMAS, and WPO alone. All bounds are estimated with a 30-point bandwidth on either side of the assignment cutoff of the program of interest (Medicine at PUC). Note that these are estimated bounds and not inference bounds.
@@ 5.2.3 On the Importance of Accounting for Strategic Behavior
Finally, to illustrate more broadly the relevance of our bounding approach, we show that naively ignoring strategic behavior in the identification of treatment effects can produce misleading results. In what follows, we compare our bounding approach to “naive estimates”, that is, estimates of
@@
We look at 302 pairs $(j,k)$ which we select as follows. First, we consider all pairs $(j,k)$ such that program $j$ is available on the platform at any point between 2004 and 2010 and at least one student lists $k$ right after $j$ in his ROL. Second, we count the number of students with score $S_{j}$ within a 30-point bandwidth of the year-specific $c_{j}$ who $P_{j}=(j,k)$. We find 327 pairs $(j,k)$ with a least 50 such students These pairs involve 137 distinct programs $j$, each of them associated with a number of $k$’s ranging from one to 18. We exclude the 25 pairs $(j,k)$ associated with three of these programs $j$ because no graduates are recorded in the data for these programs. This yields a final set of 302 pairs involving 134 distinct $j$’s.
As a simple procedure to construct a sample local to each admission cutoff $c_{j}$, we first use a 30-point bandwidth on either side of that cutoff. For a number of our $j$’s of interest, there exists another program $\ell$ that uses the same placement score as $j$ and whose cutoff $c_{\ell}$ is within a 30-point distance of $c_{j}$. In these cases, we use a bandwidth smaller than 30 around the cutoff $c_{j}$ of interest, so it does not include any such $c_{\ell}$, and the bandwidth is possibly different across sides of $c_{j}$.
@@
Table 1 provides statistics on the estimated bounds we obtain for each of the two outcomes of interest. From the table, we can see that naively ignoring strategic behavior in the identification of treatment effects can produce misleading results. First, recall that the naive estimand $\mu_{j,p}$ identifies our causal parameter of interest if we assume everyone near $c_{j}$ is a truth-teller. One might expect naive estimands to always lie within our bounds since truth-telling is a special case of our behavioral assumptions. However, Table 1 shows this is not true. Under the SPO and assumptions, the naive estimates fall outside our bounds for 12% of the pairs of interest for our first outcome and 15% of the pairs of interest for our second outcome. The fact that naive estimates fall outside our bounds is consistent with strategic behavior of agents. Indeed, Part III of our SD example (Section 4.1, right after Proposition 2), shows that the distribution of the potential outcomes conditional on reported local preferences and the priority score can change sharply at the cutoff in the presence of strategic behavior. In such a case, the naive estimand $\mu_{j,p}$ no longer equals the average treatment effect and may fall outside our bounds. More broadly, this also suggests that caution is warranted in cases like the Chilean case where most but not all students submit $P$ with $|P|<K$ or very few students are assigned to their last-listed option. It may not be safe to simply assume that everyone is a truth-teller just because the cap does not bind for most people. In other words, naive estimates may be severely biased even if the share of constrained students is small.
Table 1: Summary of overall results
Graduation Graduation
from $j$ from top univ.
Total number of pairs $(j,k)$ 302 302
Number of cases with bounds identifying the TE sign 144 83
Number of cases with naive estimate outside the bounds 35 43
Share of cases with bounds identifying the TE sign 0.48 0.28
Share of cases with naive estimate outside the bounds 0.12 0.15
@@
The table summarizes the results of the implementation of our method for 302 pairs $(j,k)$ under the SPO assumption. For each of the two outcomes of interest, the table shows the numbers of pairs $(j,k)$ for which our estimated bounds identify the sign of the treatment effect of assignment to $j$ for those whose next-preferred option is $k$. It also shows the frequency at which naive RD estimates fall outside our bounds.
@@ 6 Conclusion
Centralized mechanisms for the assignment of students to educational programs are growing in popularity across the world. These systems provide a valuable source of exogenous variation through the discontinuities that they generate around admission cutoffs. This along with individual-level data on admissions, preferences, and future outcomes allows researchers to identify a wide range of causal effects of education on outcomes. The variation is useful for identification as long as we control for students’ true preferences; however, true preferences generally differ from what students report to the assignment system given that most real-world implementations of mechanisms generate incentives for students to behave strategically.
@@ A.1 Parameter of Interest
As detailed in Section V.B of Artemov et al. (2023), the literature offers two main approaches to counterfactual analysis. The first takes the submitted rank-ordered lists (ROLs), i.e., $P$, as inputs. This approach assumes that the submitted ROLs under the current policy reflect applicants’ true preferences and that individuals will submit the same lists under the counterfactual policy. However, Artemov et al. (2023) show the existence of robust equilibria in which a non-negligible fraction of participants do not submit their true preferences, even in a strategy-proof environment. [^17] Motivated by empirical evidence of non-truthful reporting that often has limited payoff consequences, Artemov et al. (2023) develop a novel solution concept—robust equilibrium—which relaxes the Bayesian Nash equilibrium by allowing for deviations from truthful reporting when the associated payoff losses are small. They argue that this finding raises important concerns about empirical methods that treat submitted ROLs as truthful representations of preferences—both under the current policy and in counterfactual scenarios.
The second approach, instead, assumes that individuals will report their true preference profile, $Q$ (potentially different from $P$), under the counterfactual policy. This approach requires the researcher to identify the true preference $Q$ in settings where agents behave strategically and misreport. Artemov et al. (2023) advocates for this second approach, arguing that a broad class of robust equilibrium concepts yield asymptotically stable matchings. The allocations produced by these equilibria can be well approximated by the stable matching induced under truthful reporting. In other words, this implies that for a wide range of reporting behaviors, the ultimate allocation will be stable, allowing us to consider counterfactual assignments as functions of the true preferences.
This paper examines the identification of parameters essential for counterfactual analyses following the second approach advocated in Artemov et al. (2023). Let  $(Y^{c},Q^{c},S^{c})$ denote the counterfactual distribution of outcomes, preferences, and scores in the population. [^18] For simplicity, we assume that $S_{j}=S$ for any $j$ in this section. Let  $\varphi^{c}(q,s)$ denote the program to which a student with true preference  $Q^{c}=q$ and grade  $S^{c}=s$ is assigned under the counterfactual assignment mechanism. [^19] This expression is derived under the assumption of a deterministic assignment mechanism, but it extends straightforwardly to cases involving extrinsic tie-breaking through randomization. For any function  $g\in\mathcal{G}$ of potential outcomes, we have:
@@
which requires that the joint distribution of certain model primitives—such as potential outcomes, scores, and preferences—remains unchanged in the counterfactual world. See, for example, Heckman and Vytlacil (2005), who discuss both the necessity of this assumption and its limitations for certain classes of counterfactuals. Notice that, it may be more reasonable to assume that the true preference $Q$ is invariant to changes in the counterfactual world, i.e., $Q^{c}\sim Q$, rather than assuming that the submitted preferences $P$ remain unchanged. Indeed, some counterfactual scenarios may alter students’ strategic behavior, leading to $P^{c}\not\sim P$, which could pose challenges for researchers seeking to analyze such counterfactuals using the first approach described in Artemov et al. (2023).
@@
As can be seen, the key ingredient needed to conduct this type of counterfactual analysis is knowledge of the average structural functions (ASF)  $s\mapsto\mathbb{E}[g(Y(j))\mid Q=q,S=s]$ for $j\in\{0,\ldots,J\}$ and $q\in\mathcal{Q}$. In practice, nonparametrically identifying each of these functions for all  $s\in\mathcal{S}$ is extremely challenging. However, an RD design can point or set identify differences of the ASF’s at a finite number of points, namely, $\mathbb{E}[g(Y(j))-g(Y(k))\mid Q=q,S=c_{j}]$ for certain values of $j$, $k$, and $q$. In data contexts with a large number of cutoffs, Bertanha (2020) proposes a consistent and asymptotically normal estimator for counterfactual weighted averages of these differences. Outside of these contexts, researchers may rely on parametric or semi-parametric assumptions to extrapolate information from finitely many points to the entire domain.
It is important to note that the continuity assumptions utilized by RD can point or set identify ASF’s for finitely many values of $j$ and $q$ but for infinitely many values of $s$ in certain subsets of $\mathcal{S}$. For example, in the SD example of the main text, $\mathbb{E}[g(Y(0))\mid Q=q,S=s]$ is identified for any $q$ and $s\leq c_{1}$. Researchers may then combine all the identifying information with assumptions on the ASF’s to construct bounds on $\mathbb{E}[g(Y^{c})]-\mathbb{E}[g(Y)].$ Following the ideas proposed by Mogstad et al. (2018), one may assume, for example, a Bernstein polynomial basis representation for each ASF:
$\mathbb{E}[g(Y(j))\mid Q=q,S=s]=\sum_{k=0}^{K}\theta_{jk}^{q}b_{k}(s),$
where  $\{b_{k}\}_{k=0}^{K}$ are known basis functions of $s$, and  $\theta_{j}^{q}:=(\theta_{j0}^{q},\ldots,\theta_{jK}^{q})$ are unknown parameters to be (set) identified.
Finally, due to data constraints inherent to the local nature of our RD estimation, our parameter of interest in the main text does not condition on the full preference profile  $Q=q$. Instead, we follow the intuition of Kirkeboen et al. (2016) and condition on what we define in Section 3 as the local preference  $Q_{j}$. One value of $Q_{j}$ maps to multiple values of $Q$ for individuals with $S=c_{j}$. It follows that our ASF parameters in the main text
@@
As discussed earlier, parameters of this form are essential for implementing a broad class of counterfactual analyses. Their identification is generally more challenging than that of parameters conditioned on $P$, such as $\mathbb{E}\left[g(Y(j))\,\middle|\,P_{j}=(j,k),\;S_{j}=c_{j}\right].$ We do not consider these parameters because they are not directly useful to the counterfactual approach advocated by Artemov et al. (2023) for the reasons given above.
@@ A.2 Partial Orders Dominate Nonpartial Orders
First, we state some definitions. The strategy for each student $\omega$ is an ordered list of schools $P(\omega)\subseteq\mathcal{J}$ that has at least one and at most $K<J$ schools in it. The strategy profile of the economy is a correspondence $P:\Omega\rightrightarrows\mathcal{J}$ (random set). The score profile of individuals in the economy is denoted by the random vector $\bm{\mathrm{S}}:\Omega\rightarrow\mathcal{S}$. A mechanism $\varphi$ takes the whole correspondence $P$ and function $\bm{\mathrm{S}}$ as givens and produces school assignments for each individual $\omega\in\Omega$, such that $\varphi(P,\bm{\mathrm{S}}):\Omega\to\mathcal{J}^{0}$. The assignment of student $\omega$ for profiles $(P,\bm{\mathrm{S}})$ is $\varphi(P,\bm{\mathrm{S}})[\omega]$.
@@
This proof is from Haeringer and Klijn (2008), Lemma 4.2. We expand it here in terms of our framework and definitions of partial order.
@@
Given that $p^{\prime}$ is a weak partial order of both $Q(\omega)$ and $q$, for any two options $d,d^{\prime}$ in $p^{\prime}$, we have $d^{\prime}~\bar{q}~d$ implies $d^{\prime}\bar{Q}(\omega)d$. Therefore,
@@
It follows that any strong partial order $p^{\prime\prime}$ of $Q(\omega)$ that contains the same acceptable schools as $p^{\prime}$ weakly dominates $p^{\prime}$. To see this, repeat the argument above by replacing $p$ with $p^{\prime}$ and replacing $p^{\prime}$ with $p^{\prime\prime}$.

### v4

@@ 5.2.2 Treatment Effects
Figure 5 presents the estimated bounds. The left panel shows treatment effects on the probability of graduating from Medicine at PUC Santiago; the right panel shows treatment effects on the probability of graduating from a top university. Starting with the left panel, we find, reassuringly, that assignment to Medicine at PUC increases the probability of graduating from that program. In all five cases, bounds derived under any of the four sets of assumptions exclude zero, identifying a positive treatment effect. Importantly, the magnitude of the effect varies across next-best alternatives, as reflected by non-overlapping bounds. For example, when it comes to graduating from Medicine at PUC, students whose second-best is Science at PUC appear to benefit less from admission to Medicine at PUC than those whose next-best is Medicine at UChile or at Universidad de Santiago de Chile.
The right panel reveals similar patterns. Admission to Medicine at PUC increases the probability of graduating from a top university in many cases, and regardless of whether the next-best option is at a top university itself. For instance, the estimated bounds indicate a positive effect both for students whose next-best program is Science at PUC and for those whose next-best is Medicine at Universidad de Concepción or Universidad de Santiago de Chile. There again, the bounds highlight substantial heterogeneity across next-best alternatives.
As we relax assumptions, the bounds widen and become less informative. This exercise thus helps clarify the identifying power of each assumption, providing researchers with a framework to evaluate the empirical content of different preference assumptions in similar settings.
@@
The figure shows estimated bounds for the average effect on the probability of graduation (y-axis) of assigning students to Medicine at PUC as opposed to assigning them to a second-best program. Bounds are shown for five different second-best programs (x-axis), two binary outcomes (graduation from Medicine at PUC on the left panel and graduation from a top university, i.e., PUC or UChile, on the right panel), and four behavioral assumptions used to construct $\bm{\mathrm{Q}}_{j}$ (four bars for each second-best program). The four behavioral assumptions are, from left to right, SPO+UMAS+fields, SPO+UMAS, WPO+UMAS+fields, and WPO+UMAS. All bounds are estimated with a 30-point bandwidth on either side of the assignment cutoff of the program of interest (Medicine at PUC). Note that these are estimated bounds and not inference bounds.
@@ 5.2.3 On the Importance of Accounting for Strategic Behavior
Finally, we illustrate our bounding approach more broadly using 302 pairs $(j,k)$ of interest. We show that ignoring strategic behavior may be misleading when interest lies in a average treatment effect conditional on true preferences as in (3). In what follows, we compare our estimated bounds to point estimates produced by an RD that controls for reported preferences, as the current practice advocates. Current-practice estimates are consistent for
@@
Under the assumption that every student is a truth-teller at the cutoff $c_{j}$, the parameter $\mu_{j,p}$ coincides with $\mathbb{E}[Y(j)-Y(k)\mid Q_{j}=(j,k),S_{j}=c_{j}]$, which is our parameter of interest (3). Since the truth-telling assumption implies our behavioral assumptions, our bounds contain $\mu_{j,p}$ under truth-telling. On the contrary, under strategic behavior, $\mu_{j,p}$ does not equal $\mathbb{E}[Y(j)-Y(k)\mid Q_{j}=(j,k),S_{j}=c_{j}]$ in general, and our bounds may or may not contain $\mu_{j,p}$. The interpretation of $\mu_{j,p}$ depends on the type of strategic behavior. First, as mentioned earlier, strategic behavior may be of the discontinuous type to the point of breaking the continuity of $\mathbb{E}[Y(d)\mid P_{j}=(j,k),S_{j}=s]$ as a function of $s$ at $s=c_{j}$, for $d=j$ or $d=k$ (Figure 3 presents evidence consistent with this behavior). In this case, the RD is not internally valid and $\mu_{j,p}$ does not equal an average treatment effect. Alternatively, strategic behavior may not be of the discontinuous type but still make $P_{j}\neq Q_{j}$. In this case, $\mu_{j,p}$ equals a weighted average of various average treatment effect parameters, where the weights are functions of the unknown mapping $\omega\mapsto Q_{j}(\omega)$.
@@
We look at 302 pairs $(j,k)$ which we select as follows. First, we consider all pairs $(j,k)$ such that program $j$ is available on the platform at any point between 2004 and 2010 and at least one student lists $k$ right after $j$ in his ROL. Second, we count the number of students with score $S_{j}$ within a 30-point bandwidth of the year-specific $c_{j}$ who $P_{j}=(j,k)$. [^20] For a number of $j$’s of interest, there exists another program $\ell$ that uses the same placement score as $j$ and whose cutoff $c_{\ell}$ is within a 30-point distance of $c_{j}$. In these cases, we use a bandwidth smaller than 30 around the cutoff $c_{j}$ of interest, so the sample of $c_{j}$ does not include observations from the sample of $c_{\ell}$ that are subject to a different treatment change. We find 327 pairs $(j,k)$ with a least 50 such students. These pairs involve 137 distinct programs $j$, each of them associated with a number of $k$’s ranging from one to 18. We exclude the 25 pairs $(j,k)$ associated with three of these programs $j$ because no graduates are recorded in the data for these programs. This yields a final set of 302 pairs involving 134 distinct $j$’s.
@@
Table 1 provides statistics on the estimated bounds that we obtain for our two outcomes of interest, both under the SPO and the SPO+UMAS+fields assumptions. First, we see that SPO bounds identify the sign of our average treatment effect parameter for many pairs; the estimated bounds even collapse to point-estimates for a small share of pairs. Adding the UMAS and fields assumptions make our estimated bounds narrower, and they identify the sign of average effects for 70% of pairs when outcome is graduation from $j$; they collapse to point estimates for 23% of pairs when outcome is graduation from top university. Most importantly, we see that current-practice estimates fall outside of bounds in a significant share of cases, 12% & 15% under SPO and 25% & 24% under SPO+UMAS+fields. This evidence is consistent with lack of truth-telling at the cutoff of these pairs. More broadly, this also suggests caution in contexts like the Chilean case, where most but not all students submit $P$ with $|P|<K$ or very few students are assigned to their last-listed option. It may not be safe to simply assume that everyone is a truth-teller just because the constraint does not bind for most people. In other words, current-practice estimates may be severely biased for $\mathbb{E}[Y(j)-Y(k)\mid Q_{j}=(j,k),S_{j}=c_{j}]$ even if the share of constrained students is small.
Table 1: Bounds at 302 pairs $(j,k)$
Grad. from $j$ Grad. from top uni
SPO SPO+UMAS SPO SPO+UMAS
+fields +fields
Share of pairs for which:
- bounds identify sign of parameter of interest 0.48 0.70 0.28 0.38
- current-practice estimate is outside of bounds 0.12 0.25 0.15 0.24
- bounds point-identify parameter of interest 0.01 0.03 0.17 0.23
@@
The table summarizes the results of the implementation of our method for 302 pairs $(j,k)$ under the SPO and SPO+UMAS+fields assumptions. For each of the two outcomes of interest, the table shows the share of pairs $(j,k)$ for which our estimated bounds (i) identify the sign of the average treatment effect parameter (3), (ii) do not contain current-practice estimate $\widehat{\mu}_{j,p}$, and (iii) collapse to a point estimate.
@@ A.1 Parameter of Interest
As detailed in Section V.B of Artemov2023, the literature offers two main approaches to counterfactual analysis. The first takes the submitted rank-ordered lists (ROLs), i.e., $P$, as inputs. This approach assumes that the submitted ROLs under the current policy reflect applicants’ true preferences and that individuals will submit the same lists under the counterfactual policy. However, Artemov2023 show the existence of robust equilibria in which a non-negligible fraction of participants do not submit their true preferences, even in a strategy-proof environment. [^21] Motivated by empirical evidence of non-truthful reporting that often has limited payoff consequences, Artemov2023 develop a novel solution concept—robust equilibrium—which relaxes the Bayesian Nash equilibrium by allowing for deviations from truthful reporting when the associated payoff losses are small. They argue that this finding raises important concerns about empirical methods that treat submitted ROLs as truthful representations of preferences—both under the current policy and in counterfactual scenarios.
The second approach, instead, assumes that individuals will report their true preference profile, $Q$ (potentially different from $P$), under the counterfactual policy. This approach requires the researcher to identify the true preference $Q$ in settings where agents behave strategically and misreport. Artemov2023 advocates for this second approach, arguing that a broad class of robust equilibrium concepts yield asymptotically stable matchings. The allocations produced by these equilibria can be well approximated by the stable matching induced under truthful reporting. In other words, this implies that for a wide range of reporting behaviors, the ultimate allocation will be stable, allowing us to consider counterfactual assignments as functions of the true preferences.
This paper examines the identification of parameters essential for counterfactual analyses following the second approach advocated in Artemov2023. Let  $(Y^{c},Q^{c},S^{c})$ denote the counterfactual distribution of outcomes, preferences, and scores in the population. [^22] For simplicity, we assume that $S_{j}=S$ for any $j$ in this section. Let  $\varphi^{c}(q,s)$ denote the program to which a student with true preference  $Q^{c}=q$ and grade  $S^{c}=s$ is assigned under the counterfactual assignment mechanism. [^23] This expression is derived under the assumption of a deterministic assignment mechanism, but it extends straightforwardly to cases involving extrinsic tie-breaking through randomization. For any function  $g\in\mathcal{G}$ of potential outcomes, we have:
@@
which requires that the joint distribution of certain model primitives—such as potential outcomes, scores, and preferences—remains unchanged in the counterfactual world. See, for example, heckman_vytlacil_2005, who discuss both the necessity of this assumption and its limitations for certain classes of counterfactuals. Notice that, it may be more reasonable to assume that the true preference $Q$ is invariant to changes in the counterfactual world, i.e., $Q^{c}\sim Q$, rather than assuming that the submitted preferences $P$ remain unchanged. Indeed, some counterfactual scenarios may alter students’ strategic behavior, leading to $P^{c}\not\sim P$, which could pose challenges for researchers seeking to analyze such counterfactuals using the first approach described in Artemov2023.
@@
As can be seen, the key ingredient needed to conduct this type of counterfactual analysis is knowledge of the average structural functions (ASF)  $s\mapsto\mathbb{E}[g(Y(j))\mid Q=q,S=s]$ for $j\in\{0,\ldots,J\}$ and $q\in\mathcal{Q}$. In practice, nonparametrically identifying each of these functions for all  $s\in\mathcal{S}$ is extremely challenging. However, an RD design can point or set identify differences of the ASF’s at a finite number of points, namely, $\mathbb{E}[g(Y(j))-g(Y(k))\mid Q=q,S=c_{j}]$ for certain values of $j$, $k$, and $q$. In data contexts with a large number of cutoffs, bertanha_joe2020 proposes a consistent and asymptotically normal estimator for counterfactual weighted averages of these differences. Outside of these contexts, researchers may rely on parametric or semi-parametric assumptions to extrapolate information from finitely many points to the entire domain.
It is important to note that the continuity assumptions utilized by RD can point or set identify ASF’s for finitely many values of $j$ and $q$ but for infinitely many values of $s$ in certain subsets of $\mathcal{S}$. For example, in the SD example of the main text, $\mathbb{E}[g(Y(0))\mid Q=q,S=s]$ is identified for any $q$ and $s\leq c_{1}$. Researchers may then combine all the identifying information with assumptions on the ASF’s to construct bounds on $\mathbb{E}[g(Y^{c})]-\mathbb{E}[g(Y)].$ Following the ideas proposed by mogstad_santos_torgovitsky_2018, one may assume, for example, a Bernstein polynomial basis representation for each ASF:
$\mathbb{E}[g(Y(j))\mid Q=q,S=s]=\sum_{a=0}^{A}\theta_{ja}^{q}b_{a}(s),$
where  $\{b_{a}\}_{a=0}^{A}$ are known basis functions of $s$, and  $\theta_{j}^{q}:=(\theta_{j0}^{q},\ldots,\theta_{jA}^{q})$ are unknown parameters to be (set) identified.
Finally, due to data constraints inherent to the local nature of our RD estimation, our parameter of interest in the main text does not condition on the full preference profile  $Q=q$. Instead, we follow the intuition of kirkeboen2016 and condition on what we define in Section 3 as the local preference  $Q_{j}$. One value of $Q_{j}$ maps to multiple values of $Q$ for individuals with $S=c_{j}$. It follows that our ASF parameters in the main text
@@
As discussed earlier, parameters of this form are essential for implementing a broad class of counterfactual analyses. Their identification is generally more challenging than that of parameters conditioned on $P$, such as $\mathbb{E}\left[g(Y(j))\,\middle|\,P_{j}=(j,k),\;S_{j}=c_{j}\right].$ We do not consider these parameters because they are not directly useful to the counterfactual approach advocated by Artemov2023 for the reasons given above.
@@ A.2 Partial Orders Dominate Nonpartial Orders
First, we state some definitions. The strategy for each student $\omega$ is an ordered list of schools $P(\omega)\subseteq\mathcal{J}$ that has at least one and at most $K<J$ schools in it. The strategy profile of the economy is a mapping $P$ from $\Omega$ to an ordered subset of $\mathcal{J}$. We refer to this mapping as $P:\Omega\rightrightarrows\mathcal{J}$ in a slight abuse of notation. The score profile of individuals in the economy is denoted by the random vector $\bm{\mathrm{S}}:\Omega\rightarrow\mathcal{S}$. A mechanism $\varphi$ takes the whole correspondence $P$ and function $\bm{\mathrm{S}}$ as givens and produces school assignments for each individual $\omega\in\Omega$, such that $\varphi(P,\bm{\mathrm{S}}):\Omega\to\mathcal{J}^{0}$. The assignment of student $\omega$ for profiles $(P,\bm{\mathrm{S}})$ is $\varphi(P,\bm{\mathrm{S}})[\omega]$.
@@
This proof is from haeringer2008, Lemma 4.2. We expand it here in terms of our framework and definitions of partial order.
@@
Given that $p^{\prime}$ is a weak partial order of both $Q(\omega)$ and $q$, for any two options $d,d^{\prime}$ in $p^{\prime}\cup\{0\}$, we have $d^{\prime}~\bar{q}~d$ implies $d^{\prime}\bar{Q}(\omega)d$. Therefore,
@@
It follows that any strong partial order $p^{\prime\prime}$ of $Q(\omega)$ that has all the acceptable schools from $p^{\prime}$ weakly dominates $p^{\prime}$. To see this, repeat the argument above by replacing $p$ with $p^{\prime}$ and replacing $p^{\prime}$ with $p^{\prime\prime}$.